### v1

# Two Tales of Persona in LLMs: A Survey of Role-Playing and Personalization

Abstract
Recently, methods investigating how to adapt large language models (LLMs) for specific scenarios have gained great attention. Particularly, the concept of persona, originally adopted in dialogue literature, has re-surged as a promising avenue. However, the growing research on persona is relatively disorganized, lacking a systematic overview. To close the gap, we present a comprehensive survey to categorize the current state of the field. We identify two lines of research, namely (1) LLM Role-Playing, where personas are assigned to LLMs, and (2) LLM Personalization, where LLMs take care of user personas. To the best of our knowledge, we present the first survey tailored for LLM role-playing and LLM personalization under the unified view of persona, including taxonomy, current challenges, and potential directions. To foster future endeavors, we actively maintain a paper collection available to the community. [^1] https://github.com/MiuLab/PersonaLLM-Survey **footnotetext: Equal contribution.
Two Tales of Persona in LLMs:
A Survey of Role-Playing and Personalization
Yu-Min Tseng*αβ Yu-Chao Huang*α Teng-Yun Hsiao*α Yu-Ching Hsuα Jia-Yin Fooα Chao-Wei Huangα Yun-Nung Chenα αNational Taiwan University βAcademia Sinica ymtseng@nlg.csie.ntu.edu.tw {R11222015, B10502058}@ntu.edu.tw y.v.chen@ieee.org
Figure 1: An illustrated overview of the survey. In Role-Playing, LLMs act according to assigned personas (i.e., roles) under a defined environment. For example, given role names (red) with descriptions (black), LLMs are tasked to role-play in a social simulation scenario. For Personalization, LLMs consider users’ personas (blue) to generate tailored responses to the same question. The dashed rectangles indicate prompts and the solid rectangles indicate LLMs’ responses.
1 Introduction
The striking capabilities of large language models (LLMs), exemplified by ChatGPT OpenAI (2022), have significantly advanced the field of natural language processing (NLP; Wei et al., 2023; Madaan et al., 2024; Shinn et al., 2024). Recently, in addition to using LLMs as NLP task solvers or general-purpose chatbots, the question of how to adapt LLMs for specific scenarios has received great attention. To this end, leveraging personas has resurfaced as an ideal lens for adapting LLMs in target scenarios (Chen et al., 2023a, 2024). With personas, LLMs are equipped to generate tailored responses while harnessing the full capabilities. However, the growing literature on persona in the LLM era is relatively disorganized, lacking a unifying overview.
In this paper, we aim to close the gap by offering a comprehensive survey and a systematic categorization of present studies. Specifically, we divide current research into two main directions, namely LLM Role-Playing and LLM Personalization, as illustrated in Figure 1. The primary distinction is that – in role-playing, the persona belongs to the LLM, while in personalization, the persona belongs to the user. Detailed definitions and key design aspects of each are:
  • LLM Role-Playing. LLMs are tasked to play the assigned personas (i.e., roles) and act accordance to environmental feedback. The key aspect is how LLMs adapt to defined environments.
  • LLM Personalization. LLMs are tasked to take care of users’ personas (e.g., background information, or historical behaviors) to meet customized needs. The key aspect is how LLMs adapt to distinct users.
To the best of our knowledge, we present the first survey for LLM role-playing and LLM personalization under the unified view of persona. To foster future endeavors, we actively maintain a paper collection available to the research community. We hope this work could serve as a beginner guide for newcomers and a practical roadmap for current researchers in the field.
Methodology Dataset
{forest}
for tree= forked edges, draw=edgeColor, thick, font=, fill=tnodeColor, rectangle, rounded corners=4pt, text=textColor, blur shadow=shadow scale=0.95, shadow xshift=.5ex, shadow yshift=-.5ex, shadow opacity=0.25, edge=-, draw=edgeColor, line width=2pt, grow=0, child anchor=west, parent anchor=east, anchor=west, align=center, l sep+=0.3cm, s sep+=0.3cm, , root/.style= fill=Gray!15, font=, rounded corners=6pt, text width=4cm, /tikz/align=center, inner sep=6pt, tnode2_1/.style=fill=Gray!15, font=, text width=9cm, /tikz/align=center, inner sep=6pt, tnode2_2/.style=fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=6pt, tnode2_3_1/.style=fill=Gray!15, font=, text width=5cm, /tikz/align=center, inner sep=6pt, tnode2_3_2/.style=fill=Gray!15, font=, text width=9cm, /tikz/align=center, inner sep=6pt, tnode2_4/.style=fill=Gray!15, font=, text width=8cm, /tikz/align=center, inner sep=4pt, tnode3_1/.style=fill=Gray!15, font=, text width=8cm, /tikz/align=center, inner sep=6pt, tnode3_2/.style=fill=Gray!15, font=, text width=7cm, /tikz/align=center, inner sep=6pt, tnode3_3/.style=fill=Gray!15, font=, text width=8cm, /tikz/align=center, inner sep=6pt, tnode3_4/.style=fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=6pt, tnode4_1/.style=fill=tnodeColor4, font=, text width=11cm, /tikz/align=center, inner sep=6pt, tnode4_2/.style=fill=tnodeColor4, font=, text width=8.2cm, /tikz/align=center, inner sep=6pt, cnode_1/.style= fill=Yellow!15, font=, edge=dotted, text width=16cm, align=left, cnode_2/.style= fill=Yellow!15, font=, edge=dotted, text width=13cm, align=left, cnode2_1/.style=fill=Green!15, font=, edge=dotted, text width=16cm, align=left, cnode2_2/.style=fill=Green!15, font=, edge=dotted, text width=13cm, align=left, cnode3/.style=fill=tnodeColor4, font=, edge=dotted, text width=22cm, align=left, [Taxonomy, root [LLM Personalization Section 3, tnode2_1 [Methods Section 3.2, tnode2_2 [Prompting, tnode2_3_1 [Profile-Augmented, tnode2_4 [ Richardson et al. (2023);
 ONCE (Liu et al., 2023b), fill=Yellow!15, font=, edge=dotted, text width=11cm, align=left] ] [Retrieval-Augmented, tnode2_4 [ WorkSM (Mysore et al., 2023);
 LAMP (Salemi et al., 2024), fill=Green!15, font=, edge=dotted, text width=11cm, align=left] [ Li et al. (2023a);
 PEARL (Mysore et al., 2023), fill=Yellow!15, font=, edge=dotted, text width=11cm, align=left] ] [Vanilla, tnode2_4 [ Dai et al. (2023);
 Christakopoulou et al. (2023);
 BookGPT (Zhiyuli et al., 2023), fill=Yellow!15, font=, edge=dotted, text width=11cm, align=left] ] ] [Retrieval Augmentation, tnode2_3_2 [ PER-CHAT (Wu et al., 2021);
 PersonaKGC (Fu et al., 2022);
 DuLemon (Xu et al., 2022);
 PerLTQA (Du et al., 2024), cnode2_1] [ Wu et al. (2021);  Fu et al. (2022);
 PLATO-LTM (Xu et al., 2022);  Du et al. (2024);
 MemoryBank (Zhong et al., 2024), cnode_1] ] [Fine-Tuning, tnode2_3_2 [ P-RLHF (Li et al., 2024c), cnode_1] ] ] [Tasks Section 3.1, tnode2_2 [Personalized Education, tnode2_3_2 [ HumSum Shehata et al. (2023);
 EduChat (Dan et al., 2023);  Park et al. (2024), cnode_1] ] [Personalized Healthcare, tnode2_3_2 [ MedDialog (Zeng et al., 2020);
 IMCS-21 (Chen et al., 2022b);
 iCliniq (Li et al., 2023f), cnode2_1] [ openCHA (Abbasian et al., 2024a);
 CHA (Abbasian et al., 2024c);
 MaLP (Zhang et al., 2024a);
 HealthLLM (Jin et al., 2024b), cnode_1], ] [Personalized Search, tnode2_3_2 [ Pass et al. (2006), cnode2_1] [ CoPS (Zhou et al., 2024), cnode_1] ] [Recommendation System, tnode2_3_2 [ MovieLens (Harper and Konstan, 2015), cnode2_1] [ PALR (Yang et al., 2023b);  Dai et al. (2023);
 Christakopoulou et al. (2023);
 BookGPT (Zhiyuli et al., 2023);
 ONCE (Liu et al., 2023b), cnode_1] ] [Personalized
Dialogue, fill=Gray!15, font=
, text width=5cm, /tikz/align=center, inner sep=6pt [User Persona
Modeling, fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=4pt [ PersonaChat (Zhang et al., 2018a);
 ConvAI2 PersonaChat (Dinan et al., 2019), cnode2_2] [ $\mathcal{P}^{2}$ (Liu et al., 2020);  Kim et al. (2024b);
 CLV (Tang et al., 2023b), cnode_2] ] [ToD Modeling, fill=Gray!15, font=, text width=6cm, /tikz/align=center, inner sep=4pt [ MutiWoz (Budzianowski et al., 2018);
 SGD (Rastogi et al., 2020), cnode2_2] [ RefGPT (Yang et al., 2023a);
 ProToD (Hu et al., 2023);
 DSP (Li et al., 2024e), cnode_2] ] ] ] ] [LLM Role-Playing Section 2, tnode3_1 [Agentic
Interactions Section 2.3, tnode3_2 [Emergent Behaviors, fill=Gray!15, font=, text width=7cm, /tikz/align=center, inner sep=6pt [ AgentVerse (Chen et al., 2023c);  Fu et al. (2023);
 MedAgents (Tang et al., 2023a);  CAMEL (Li et al., 2024b), fill=Yellow!15, font=
, edge=dotted, text width=18cm, align=left] ] [Schemas, fill=Gray!15, font=, text width=5cm, /tikz/align=center, inner sep=6pt [Multi-Agent, tnode3_4 [ ChatDev (Qian et al., 2023);
 MedAgents (Tang et al., 2023a), fill=Yellow!15, font=
, edge=dotted, text width=12.8cm, align=left] ] [Single-Agent, tnode3_4 [ Voyager (Wang et al., 2023a);
 MindAct (Deng et al., 2024), fill=Yellow!15, font=, edge=dotted, text width=12.8cm, align=left] ] ] ] [Environments Section 2.2, tnode3_2 [General Framework, tnode3_3 [ CAMEL AI Society; CAMEL Code; CAMEL Math;
 CAMEL Science (Li et al., 2024b), fill=Green!15, font=, edge=dotted, text width=17cm, align=left] [ CAMEL (Li et al., 2024b);  SPP (Wang et al., 2023e);
 AgentVerse (Chen et al., 2023c), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] [LLM as Evaluators, tnode3_3 [ DRPE (Wu et al., 2023b);  ChatEval (Chan et al., 2023), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] [Medical Application, tnode3_3 [ Wu et al. (2023a);  MedAgents (Tang et al., 2023a), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] [Game, fill=Gray!15, font=
, text width=5cm, /tikz/align=center, inner sep=6pt [ Generative Agents (Park et al., 2023);
 Humanoid Agent (Wang et al., 2023f);
 Fu et al. (2023);  Voyager (Wang et al., 2023a), fill=Yellow!15, font=
, edge=dotted, text width=20cm, align=left] ] [Web, fill=Gray!15, font=, text width=5cm, /tikz/align=center, inner sep=6pt [ Mind2Web (Deng et al., 2024);  WebArena (Zhou et al., 2023);
 VisualWebArena (Koh et al., 2024);
 VisualWebBench (Liu et al., 2024), fill=Green!15, font=
, edge=dotted, text width=20cm, align=left] [ MindAct (Deng et al., 2024);  CogAgent (Hong et al., 2023b);
 SeeClick (Cheng et al., 2024a), fill=Yellow!15, font=, edge=dotted, text width=20cm, align=left] ] [Software Development, tnode3_3 [ SoftwareDev (Hong et al., 2023a);
 SRDD (Qian et al., 2023), fill=Green!15, font=, edge=dotted, text width=17cm, align=left] [ Dong et al. (2023);  ChatDev (Qian et al., 2023);
 MetaGPT (Hong et al., 2023a), fill=Yellow!15, font=, edge=dotted, text width=17cm, align=left] ] ] ] ]
Figure 2: The taxonomy of LLM role-playing and LLM personalization.
Overview.
We organize the survey paper as laid out in the taxonomy (Figure 2). Firstly, we introduce LLM role-playing Section 2, including agent-based role-playing of LLMs Section 2.1, various environments LLMs adapt to Section 2.2, and how LLMs interact within the environments as well as emergent behaviors stemmed from their interactions Section 2.3. Subsequently, we elaborate LLM personalization Section 3, outlining the diverse tasks Section 3.1 and methods Section 3.2 for personalizing LLMs. The benchmarks and datasets for various environments and personalized scenarios are introduced in relevant sections accordingly. We also add a separate branch Section 4 for LLM personality evaluation focusing on psychological assessments. Lastly, we highlight current challenges in role-playing and personalized LLMs and suggest potential future directions Section 5.
2 LLM Role-Playing
Figure 3: An illustration of five LLM role-playing environments: Software Development Section 2.2.1, Web Section 2.2.2, Game Section 2.2.3, Medical Application Section 2.2.4, and LLM as Evaluators Section 2.2.5. For each environment, we provide a simple scenario with a task description (red-bordered) and relevant personas (i.e., roles; blue-bordered). The dashed rectangle represents an example LLM role-playing prompt template. In addition to the above environments, several research also proposes general frameworks applicable to different environments Section 2.2.6.
In order to perform diverse tasks effectively, a potential solution is to train specialized LLMs tailored to each task. However, this process is computationally prohibitive and time-consuming. To address this, recent studies have demonstrated that leveraging role-playing increases LLMs capabilities, including reasoning (Wu et al., 2023a; Kong et al., 2023), planning Huang et al. (2024); Li et al. (2024b), and problem-solving (Wang et al., 2023e; Qian et al., 2023). Specifically, role-playing triggers the corresponding inherent personalities of LLMs to generate responses aligned with given roles and environments. Equipped with assigned personas (i.e., roles), LLMs adapt adequately to defined environments, thereby enhancing their performance.
In this section, we first clarify the relationship between LLM role-playing and LLM-based agent  Section 2.1. Next, we introduce diverse environments that LLMs could adapt to Section 2.2. Lastly, we identify agentic interactions between LLMs within defined environments and emergent behaviors stemming from their interactions Section 2.3.
2.1 LLM-based Agent
Recently, several research has explored the planning, reflection, and tool-using abilities in LLMs, potentially paving the way for artificial general intelligence (AGI; Bubeck et al., 2023). Thus, LLM-based agent has surged as an fast-growing research field (Wang et al., 2024b; Xi et al., 2023). Moreover, the Multi-Agent framework, where multiple LLM-based agents cooperate and communicate with each other in order to solve complex tasks, has also appeared as a prevalent research direction (Guo et al., 2024).
To answer the key question discussed in Section 1: how LLMs adapt to defined environments, the predominant approach is LLM-based agents with role-playing, where LLMs act according to the assigned personas to interact with the environments. For instance, one of the first works of LLM-based agents proposes generative agents (Park et al., 2023), which adopts role-playing to mimic human behaviors. By giving a list of sentences with name, age, and personality trait that specifies the role, generative agents can perform accordingly and engage in a social simulation environment.
2.2 Environments
One key aspect of the agentic frameworks is the environments, within which the LLMs could operate and interact given certain rules and receive feedback. We classify the environments for LLM role-playing into five categories: software development, web, game, medical application, and LLM as evaluators, detailed in Section 2.2.1 to Section 2.2.5. Furthermore, we explore a range of frameworks that generalize across different environments in Section 2.2.6.  Figure 3 demonstrates an illustration by providing a simple example task of each environment.
2.2.1 Software Development
The tasks of software development are to design programs, software, or coding projects. Example tasks include “Create a snake game.”, “Create a Python program to develop an interactive weather dashboard.” (Hong et al., 2023a), and HumanEval (Chen et al., 2021). Due to the complexity of these tasks, often too intricate to be completed correctly on the first attempt, existing research predominantly leverages the Waterfall model Petersen et al. (2009); Bassil (2012) or Standardized Operating Procedures (SOPs) Belbin and Brown (2022); DeMarco and Lister (2013) to break down the tasks into manageable sub-tasks.
Similar to real-world software development projects, a collaborative multi-agent framework with LLM role-playing operating as a company serves as an effective approach for software development environment Qian et al. (2023); Hong et al. (2023a); Dong et al. (2023). Concretely, the roles might include Chief Technology Officer (CTO), Chief Product Officer (CPO), Chief Executive Officer (CEO), Product Manager, Engineer, Reviewer, and Tester. By assigning specific roles, LLMs are capable of generating content in a step-by-step and accurate manner.
A recent work by Dong et al. (2023) proposes one of the first self-collaboration frameworks that encompasses division of labor and collaboration among multiple LLM agents, each acting as a specialized “experts” to address complex code generation tasks. Following the Waterfall model, ChatDev (Qian et al., 2023) divides the development process into a four-phase pipeline: designing, coding, testing, and documenting and proposes Chat Chain to decompose each phase into a sequence of atomic sub-tasks. Differing from the above works, MetaGPT (Hong et al., 2023a) requires LLM agents to generate structured outputs instead of free-text, which demonstrates a significant increase in the success rate of target code generation.
Benchmark #Instances #Domains Realistic Dynamic Visual Assessment
Env. Interaction Needed
WebShop (Yao et al., 2022) 12,087 1 End-to-end
Mind2Web (Deng et al., 2024) 2,350 5 End-to-end
WebArena (Zhou et al., 2023) 812 4 End-to-end
VisualWebArena (Koh et al., 2024) 910 3 End-to-end
VisualWebBench (Liu et al., 2024) 1,500 12 Fine-grained
Table 1: Comparison between recent benchmarks in the web environment. Realistic Env. denotes whether the benchmark’s environments are based on actual web pages or realistic web navigation simulations. Dynamic Interaction indicates whether the benchmark supports dynamic interactions rather than remaining in static states. Visual Needed denotes whether the benchmark involves visually grounded tasks. Assessment refers to the types of assessment. An end-to-end benchmark includes tasks with simple instructions, requiring step-by-step solutions to reach the final answers. A fine-grained benchmark contains tasks with a detailed assessment of essential skills in the web environment such as Optical Character Recognition (OCR), and semantic understanding.
2.2.2 Web
In this environment, LLMs operate web navigation autonomously, performing actions such as clicking items, capturing contents, and searching from external knowledge on the web, without a specific persona assigned. Certainly, web tasks involve two key components: HTML understanding and visual grounding, which are highly related to the effectiveness of web agents (Zheng et al., 2024; Koh et al., 2024). Meanwhile, a stream of works, compiled in Table 1, proposes several benchmarks to assess web agents in diverse aspects.
HTML Understanding.
Kim et al. (2024a) showcase that the ability of HTML understanding is inherent in LLMs with the Recursive Criticism and Improvement (RCI) prompting method. However, due to the special formats and long context elements of HTML which are hard for LLMs to process and respond accurately, most research enhances this capability through fine-tuning methods (Gur et al., 2022, 2023; Deng et al., 2024).
Visual Grounding.
Another line of research focuses on the visual grounding aspect of HTML understanding, which directly operates on rendered webpages instead of the HTML source code. Some literature proposes web agent frameworks, such as CogAgent (Hong et al., 2023b) and SeeClick (Cheng et al., 2024a), leveraging Large Multi-modal Models (LMMs) (Achiam et al., 2023; Team et al., 2023). With additional information from webpage screenshots, LMMs usually outperform text-based LLMs Zheng et al. (2024).
2.2.3 Game
Prior to the rise of LLMs, foundation works in agent-based game playing are AlphaGo (Silver et al., 2016) and AlphaGo Zero (Silver et al., 2017). Additionally, Deal or No Deal? (Lewis et al., 2017), which involved bargaining between two agents based on Recurrent Neural Network (RNN) models, also demonstrated the potential capabilities of agents in game playing. More recently, LLMs have surged as a more advanced technique for agents in a variety of game environments, including Minecraft (Wang et al., 2023a), social simulation (Park et al., 2023; Wang et al., 2023f), and bargaining game (Fu et al., 2023).
Within these environments, LLMs are able to role-play as a general assistant Wang et al. (2023a), or characters related to the environment, such as buyers and sellers Fu et al. (2023). Especially, gaming environments usually contain a wide range of information, including settings, utilizable tools, and nearby situations, which is hard for LLMs to memorize and response. Thus, memory stream with retrieval-based approaches are a crucial component for the effectiveness of LLM-based agents’ role-play in the game environments (Park et al., 2023; Wang et al., 2023a).
2.2.4 Medical Application
Prior works have shown the potential utility of LLMs in medicine, establishing the stepping stone of unleashing the domain-specific ability of LLMs in this environment (Singhal et al., 2023; Thirunavukarasu et al., 2023). Wu et al. (2023a) propose DR-CoT as the first approach of leveraging LLM role-playing on knowledge-intensive, multi-step diagnostic reasoning, exhibited a striking improvement from standard prompting. The following work, MedAgent (Tang et al., 2023a), introduces a multi-agent collaboration framework into medical reasoning through five processes: expert gathering, analysis proposition, report summarization, collaborative consultation, and decision making, to mimic actual medical scenarios.
Both studies assign medically relevant personas to LLMs, ranging from general roles like Doctor and Patient to specific ones such as Neurology Expert and Psychiatry Expert. Their research demonstrates that LLMs inherently possess medical knowledge (Liévin et al., 2024), thus enhancing the performance through LLM role-playing successfully. Nonetheless, there is still ample room for further exploration of LLM role-playing in a medical application environment.
2.2.5 LLM as Evaluators
As LLMs grow increasingly powerful, the evaluation of generated content becomes an extremely vital and challenging issue. Traditional n-gram metrics and model-based evaluations, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang et al., 2019), cannot appropriately assess alignment with human intent (Wang et al., 2023b; Chang et al., 2023b). For instance, a summary with a high ROUGE score might not be a fluent and meaningful passage, potentially misrepresenting the original article.
Therefore, the concept of LLM as evaluators (i.e., LLM-as-a-judge) has been introduced as an alternative way of evaluation. Specifically, LLMs are capable of assessing human-like values of model responses, such as empathy, fluency, and reasonableness. Several studies also demonstrate that the judgments made by LLMs have a higher correlation with human ground truth than traditional metrics (Chiang and Lee, 2023; Zheng et al., 2023; Wang et al., 2023d; Lin and Chen, 2023).
Aiming for greater similarity with human evaluation, roles in this environment span a broad spectrum representing various perspectives of human beings in society, such as the general public, critic, and news author.  Wu et al. (2023b) propose the DRPE method to assess the quality of summarization by assigning statically objective roles, and dynamically subjective roles based on task settings to LLMs. Another work, ChatEval (Chan et al., 2023), further adds discussion rounds within roles to improve the evaluation process, simulating a judge group in reality. In this way, LLM as evaluators could serve as an alternative solution of evaluation while models become stronger.
2.2.6 General Framework
Despite the effectiveness of the aforementioned LLM role-playing frameworks tailored to each task, a line of research indicates that these personas heavily rely on human assignment, requiring prior knowledge and understanding of the tasks (Chen et al., 2023c). Consequently, enhancing the generalizability of these frameworks and employing automatic prompt crafting appear as potential solutions to these limitations (Li et al., 2024b; Wang et al., 2023e).
For example, Li et al. (2024b) propose a novel communicative agent framework that allows LLM agents to collaborate autonomously by Inception Prompting, requiring minimal human intervention. Subsequently, Wang et al. (2023e) introduce Solo Performance Prompting (SPP), a method for LLMs to automatically identify personas based on given problems, and further engage in multi-turn self-collaboration with those personas. Another work, proposed by Chen et al. (2023c), is a multi-agent framework called AgentVerse that enables LLMs to dynamically adjust the personas of group members based on current progress. Overall, these frameworks demonstrate significant performance and generalizability across a range of tasks, including reasoning and generation.
2.3 Agentic Interactions
In this section, we first present two distinct schemas of LLM role-playing in Section 2.3.1, and further discuss the emergent behaviors of LLMs in Section 2.3.2.
2.3.1 Schemas
Across various environments, we categorize LLM role-playing schemas into two catogories: Single-Agent and Multi-Agent. The distinction is whether agents could achieve the given task single-handedly in the defined environments. Notably, each agent is capable of planning, reflecting, and decision-making (Wang et al., 2024b; Xi et al., 2023). Detailed definitions are as follows.
Single-Agent Schema.
We define the single agent schema as: one agent is able to achieve its goal independently without assistance from others, even though it may coexist with multiple agents in the defined environment.
Under the notion of LLM role-playing, single agent schema mainly appears in game and web environments, where LLMs pay more attention on environmental information and feedback rather than collaboration. For example, Voyager (Wang et al., 2023a) agents, playing general assistant roles, are tasked to continuously explore the defined environment, acquire diverse skills, and make novel discoveries in Minecraft. Despite the presence of multiple Voyager agents in Minecraft, each agent is capable of exploring the gaming world on its own.
Multi-Agent Schema.
In comparison, we define the multi-agent schema as one agent necessitated to collaborate, communicate, and ultimately achieve its goal with the support of other agents.
Software development, medical applications, and LLM as evaluators are primary environments for this schema. Similar to the real world, the essence of these environments is human interaction. LLMs impersonate various critical roles and their functionalities by adopting relevant personas, subsequently forming a multi-agent group to tackle given tasks within the defined environments. For example, AgentVerse (Chen et al., 2023c) and ChatDev (Qian et al., 2023) both propose multi-agent frameworks that exchange information and cooperate to accomplish their tasks efficiently.
As identified in prior surveys of LLM-based agents (Xi et al., 2023; Guo et al., 2024), there are two collaboration paradigms in the multi-agent schema: Cooperative and Adversarial. The cooperative paradigm facilitates information sharing among agents, with some frameworks using message pools to store each agent’s current state and ongoing tasks (Hong et al., 2023a; Tang et al., 2023a; Chen et al., 2023c). On the other hand, the adversarial paradigm, including debate, competition, and criticism, enhances the decision-making process and seeks more advantages through adopting opposing perspectives (Chan et al., 2023; Fu et al., 2023).
Figure 4: An illustration of five types of personalized LLMs: Personalized Dialogue Section 3.1.1, Personalized Search Section 3.1.3, Recommendation System Section 3.1.2, Personalized Healthcare Section 3.1.4, and Personalized Education Section 3.1.5. For dialogue and education tasks, we illustrate multi-turn interactions between LLMs and users. For other tasks, the dashed rectangles represent prompts, and the solid rectangles represent the responses of LLMs.
2.3.2 Emergent Behaviors
Recent research has explored emergent behaviors in LLM role-playing (Chen et al., 2023c; Li et al., 2024b; Wang et al., 2023e; Tang et al., 2023a; Fu et al., 2023). Furthermore, as mentioned in Cheng et al. (2024b), LLM multi-agent collaboration can reflect human society phenomena or theories in a social psychology view, such as conformity and consensus reaching. To provide a pioneering perspective for future research, we integrate three collaborative behaviors in this section, mainly following Chen et al. (2023c) and others’ findings.
Volunteer Behavior.
This behavior primarily occurs in the multi-agent schema with the cooperative collaboration paradigm. To accomplish a team goal, LLMs proactively assist their peers or inquire if there is anything they can help with. Moreover, they contribute resources to each other, such as unallocated time and possessed materials. Through this volunteer behavior, LLMs enhance team efficiency and demonstrate cohesion and commitment within defined environments (Chen et al., 2023c; Hong et al., 2023a).
Conformity Behavior.
This behavior occurs in situations where a LLM deviates from the team goal. After receiving criticism and suggestions from others, the deviating LLM refines and adjusts its behavior or decisions to better cooperate with the team. Through this conformity behavior, LLMs align with the mutual goal and pursue improved accuracy and completeness (Tang et al., 2023a; Fu et al., 2023).
Destructive Behavior.
Occasionally, LLMs undertake various actions that lead to undesired and detrimental outcomes. For instance, LLMs may exhibit a Bad Mind that seeks to control the world (Li et al., 2024b). Furthermore, LLMs consistently display toxicity or reveal deep-seated stereotypical biases when equipping personas (Deshpande et al., 2023; Gupta et al., 2023). Such destructive behavior raises safety and bias concerns of LLM role-playing. Thus, mitigating these issues is also outlined in Section 5 as a potential future direction.
3 LLM Personalization
A crucial aspect of developing strong and capable large language models is further shaping them to meet our personalized needs. Personalized LLMs have gained significant interest in recent years, leading to the organization of various workshops Chen et al. (2023e); Deshpande et al. (2024) and competitions Yusupov and Kuratov (2018); Dinan et al. (2019). These events have introduced a wealth of innovative concepts for tailoring LLMs to individual user needs. Most LLMs use reinforcement learning from human feedback to fine-tune the model, aligning the output with human expectations and preferences. However, in the process, they infuse collective consciousness into the model because of biases in human-annotated data. To make LLMs more context-sensitive and enhance the human experience by adapting to individual consciousness, LLMs need to consider individual information, historical behavior, and user personas.
In this section, we first introduce five tasks of personalized LLMs Section 3.1. Subsequently, we discuss various methods used in personalized LLMs Section 3.2, including fine-tuning methods, retrieval augmentation, and prompting methods.
3.1 Tasks
We present various aspects of personalizing LLMs across diverse tasks, including personalized dialogue Section 3.1.1, personalized recommendation system Section 3.1.2, personalized search Section 3.1.3, personalized healthcare Section 3.1.4, and personalized education Section 3.1.5, as illustrated in Figure 4.
3.1.1 Personalized Dialogue Generation
One of the most active areas for personalized LLMs is personalized dialogue generation, which can be categorized into two distinct goals: (1) generating personalized task-oriented responses and (2) modeling users’ personas.
  1. (1)
    Task-oriented Dialogue Modeling (ToD Modeling). ToD modeling guides users in completing specific tasks through multiple interactive steps, such as hotel bookings and restaurant reservations. It consists of understanding the semantic meaning of user utterances, following up with policy planning to determine the dialogue act, and finally generating an active response in form of natural language. See an example in Section A.1.
  2. (2)
    User Persona Modeling. User persona modeling detects the user persona based on dialogue history and generates customized responses tailored for each user. See an example in Section A.2.
Background of Personalized Dialogue Generation.
ToD modeling typically consists of four modules: Natural Language Understanding (NLU), Dialogue State Tracker (DST), Dialogue Policy (DP), and Natural Language Generator (NLG). Among these, DST Jacqmin et al. (2022) plays a crucial role in modeling a multi-turn dialogue while updating the state of the conversation. The state could include user information, preferences, and goals. Mrkšić et al. (2016) propose a novel method known as Neural Belief Tracker (NBT), which features an enhanced version of update mechanisms as described in Mrkšić and Vulić (2018). This method advances representation learning by predicting and updating various aspects of the user’s requests and goals through belief tracking.
Prior to the rise of LLMs, many models focus on improving different parts of the module, such as state tracking Mrkšić et al. (2017); Rastogi et al. (2018); Wu et al. (2019b); Zhang et al. (2020), while others concentrate on policy optimization using ground-truth dialogue states Wang et al. (2020); Sun et al. (2021). Various attempts tried to combine different modules to create a fully end-to-end ToD modeling: Liu et al. (2018); Yang et al. (2021a) use reinforcement learning (RL) to combine DP and NLG. Lei et al. (2018) combine DST and NLG with a sequence-to-sequence approach. Huang et al. (2020) propose a method based on the variational autoencoder (VQ-VAE) framework Kingma and Welling (2013) and use three-stage learning, including Semantic Latent Action Learning, Action Alignment across Domains, and Domain-Specific Action Learning. Finally, the SimpleTOD Hosseini-Asl et al. (2020) model integrates different sub-tasks in a unified end-to-end manner, paving the way for fully LLM-based approaches in ToD modeling.
On the other hand, in the pre-LLM era, User persona modeling used methods like Sordoni et al. (2015) and StarSpace Wu et al. (2018) to rank the most similar utterance in the dataset and generate a candidate reply. Additionally, Miller et al. (2016) enhanced its ability by considering the dialogue history.
Category Dataset #Dialogues #Utterance #Domains
ToD Section 3.1.1 MultiWOZ 1.0 Budzianowski et al. (2018) 10,438 75,894 7
MultiWOZ 2.0 Ramadan et al. (2018) 8,438 63,841 7
MultiWOZ 2.1 Eric et al. (2020) 7,032 57,022 7
MultiWOZ 2.2 Zang et al. (2020) 10,438 71,572 7
SGD Rastogi et al. (2020) 22,825 463,284 20
STAR Mosig et al. (2020) 6,652 127,833 13
AirDialogue Wei et al. (2018) 4,000 52,000 1
UniDA He et al. (2022) 70,726 975,780 13
Persona Section 3.1.1 PersonaChat Zhang et al. (2018b) 11,907 164,356 1
ConvAI2 Dinan et al. (2019) 13,500 182,150 1
Baidu PersonaChat PapersWithCode (2020) 20,000 280,000 1
JPersonaChat Sugiyama et al. (2021) 10,000 140,000 1
JEmpatheticDialogues Sugiyama et al. (2021) 25,000 350,000 1
DailyDialog Li et al. (2017) 13,118 102,979 10
Table 2: A list of commonly used datasets for ToD and Persona Modeling.
ToD Modeling.
Madotto et al. (2020) demonstrate promising results for the few-shot ability of LLMs on each module of NLU, DST, DP, and NLG, showcasing pre-trained LLMs as a favorable end-to-end approach. Early works on creating a fully end-to-end model include SimpleTOD Hosseini-Asl et al. (2020) and SOLOIST Peng et al. (2021) These models handle all sub-tasks in an end-to-end, single-sequence prediction setting on pre-trained LLMs. Similar work UBAR Yang et al. (2021b) models the entire dialogue history, including belief states, system acts, and system responses as the input sequence. Directional Stimulus Prompting (DSP) Li et al. (2024e) trains a small policy model to generate hints and guide LLMs in completing tasks. In ProToD Hu et al. (2023), inspired by DSP (Li et al., 2024e), they propose a framework that incorporates future action hints and a policy model trained with supervised learning and reinforcement learning. Hudeček and Dusek (2023) employ in-context learning, retrieval, and state tracking with instruction fine-tuned LLM. RefGPT Yang et al. (2023a) generates multi-turn, truthful dialogues by augmenting the dialogue history with factual and reliable sources and using prompts to guide the LLM in structuring according to predefined dialogue settings.
Some models combine reinforcement learning, focusing on the policy model Wu et al. (2023c), the dialogue state tracker and natural language generator Bang et al. (2023), designing reward functions for end-to-end ToD agents Feng et al. (2023), or augmenting prompts Li et al. (2024e); Hu et al. (2023).
Other works use LLMs to generate multi-turn dialogue datasets for training language models (Yang et al., 2023a; Huryn et al., 2022; Xu et al., 2023). On the application side, personalized dialogue has also been used in procedural content generation (PCG) Ashby et al. (2023) in video games for customized dialogue generation.
Dataset Scene Task #Instances #Users #Items
Amazon Review Ni et al. (2019) Products Ratings, Reviews 233.1M 43.53M 15.17M
MovieLens Harper and Konstan (2015) Movies Ratings 100,000 1,000 1,700
Yelp Yelp (2013) Businesses Ratings & Reviews 6,990,280 1,987,897 150,346
TripAdvisor Li et al. (2023c) Hotels, Restaurants Ratings & Reviews 320,023 9,765 6,280
MIND Wu et al. (2020) News Sequence recommendation 15M 1M 160k
Table 3: A list of commonly used datasets in personalized LLMs for recommendation and search task. For the fifth column, the instances include reviews and ratings.
User Persona Modeling.
Zhang et al. (2018a) present one of the earlier works on personalized dialogue generation by training DIALOGPT on Reddit comments and achieving human-like performance in a single-turn dialogue setting. Later on, Zheng et al. (2019) concatenates word embedding, positional embedding, and attribute embedding and uses an encoder-decoder to pre-train a language model. Wu et al. (2019a) propose a variational autoencoder (VAE) to generate persona-aware responses. They also introduce three persona evaluation metrics: uRank, which measures the user’s language style; uPPL, which imitates the language style of the user; and uDistinct, which measures the diversity between different users.
More sophisticated models, such as CoBert Zhong et al. (2020), propose persona-based empathetic conversations using BERT as the encoder and a two-hop co-attention mechanism Lu et al. (2017) to refine these embeddings and identify the most relevant response given the context and persona information. They further present a persona-based empathetic conversation (PEC) dataset. Other works take different approaches. For example, Song et al. (2020) utilizes natural language inference (NLI) as an RL task with response-persona as a reward to generate persona-consistent dialogue. Liu et al. (2020) proposes $\mathcal{P}^{2}$, a mutual persona perception model, and employs supervised training and self-play fine-tuning during the training process. Tang et al. (2023b) combine sparse persona descriptions, dense persona descriptions, and dialogue history to generate personalized responses.
Datasets.
In Table 2, we organize some common datasets used for ToD Modeling and Persona Modeling. Among them, various versions of MultiWOZ Budzianowski et al. (2018); Ramadan et al. (2018); Eric et al. (2020); Zang et al. (2020) and PersonaChat Zhang et al. (2018b) are the most commonly used. Updated versions of MultiWOZ improve in several aspects: data quality, dialogue complexity, schema and ontology updates, and dataset sizes. PersonaChat contains various persona profiles, consisting of background, preferences, and personality traits. These profiles enable the modeling of coherent and contextual multi-turn diverse dialogue scenarios.
On the potential application side for persona modeling, Tu et al. (2023) matches individuals with persona-compatible virtual supporters and introduces the MBTI-S2Conv dataset, which contains conversations between characters with distinct profiles. On the side of creating synthetic persona-related datasets, Lotfi et al. (2024) distill a dataset based on PersonaChat (Zhang et al., 2018b) with Big Five personality traits labels. Han et al. (2024) propose a Korean synthetic dataset with an emphasis on a Big Five personality-generating pipeline.
Takeaway.
In this section, we survey various approaches for (1) generating personalized task-oriented responses and (2) modeling users personas. Currently, LLMs demonstrate a remarkable ability to model task-oriented dialogues and personas compared to many specialized models. However, there are potential exploits in persona modeling, as discussed in Section 2.3. For example, Jin et al. (2024a); Shah et al. (2023); Deshpande et al. (2023) show that by assigning personas to LLMs, they can potentially aid in jailbreaking or produce toxic outputs. Further discussion on the challenges is outlined in Section 5.
3.1.2 Recommendation System
Paper Scene Dataset Method Task
Li et al. (2023d) Hotel, Movies & TV, Restaurant TripAdvisor, Amazon, Yelp Embeddings, Prompting, Fine-tuning Aspect extraction, Rating Prediction
P5 Geng et al. (2022) Sports, Beauty, Toys, Yelp Amazon Ni et al. (2019), Yelp Pretraining, Prompting Rating Prediction, Sequential Recommendation, Explanation Generation, Review Generation, and Direct Recommendation
PETER Li et al. (2021) Hotel, Movies & TV, Restaurant TripAdvisor, Amazon, Yelp Transformer Rating prediction and Explanation Generation
PEPLER Li et al. (2023c) Hotel, Movies, TV and Restaurant TripAdvisor5 (Hotel), Amazon (movies& TV) and Yelp7 (restaurant) Prompting, Fine-tuning Explanation Generation
PALR Yang et al. (2023b) Movies, Beauty MovieLens-1M Harper and Konstan (2015), Amazon Beauty Ni et al. (2019) Fine-tuning, User Profile Generation, Retrieval User Profile Generation and Direct Recommendation
Chu et al. (2023) Sports, Outdoors, Beauty, Toys and Games Amazon Fine-tuning Rating Prediction, Sequential Recommendation, Direct Recommendation, Explanation Generation and Review Summarization
Liu et al. (2023a) Beauty Amazon Prompting Rating Prediction, Sequential Recommendation, Direct Recommendation, Explanation Generation and Review Summarization
Zhang et al. (2023) Video Games Amazon Instruction tuning Sequential Recommendation and Direct Recommendation
Hou et al. (2024) Movies Amazon Ni et al. (2019), MovieLens-1M Harper and Konstan (2015) Prompting Sequential Recommendation
Wang and Lim (2023) Movies MovieLens-1M Harper and Konstan (2015) Prompting Sequential Recommendation and Direct Recommendation
Chen et al. (2022a) News MIND Wu et al. (2020), Reddit Fine-tuning with weak labels Direct Recommendation
Table 4: An overview of existing research in recommendation. Following the classification of Liu et al. (2023a), we classify recommendation systems into five types: rating prediction, sequential recommendation, explanation Generation, and review generation, and direct recommendation.
In this section, we introduce the field of recommendation systems, a significant research area in personalization. Recommendation systems aim to recommend items to users that match their preferences. For example, systems may recommend various items such as books, movies, and restaurants. Moreover, we compile commonly used datasets in Table 3 and an overview of existing research in Table 4.
Existing studies explore various prompting methods for using LLMs in recommendation systems. Li et al. (2023c, 2021) develop a method to efficiently incorporate a user’s personal information into pre-trained Transformer models. By using user and item ID information, they construct a recommendation system that outperforms others on the TripAdvisor, Amazon, and Yelp datasets. Li et al. (2023d) use LLMs for aspect extraction through prompt tuning. They combine aspect extraction with aspect-based recommendations. Chen et al. (2022a) propose a method to generate personalized chit-chat to enhance recommendation systems. Wang and Lim (2023) utilize a three-step prompting process to achieve better performance.
Additionally, existing studies explore different frameworks for recommendation systems. Geng et al. (2022) unify various recommendation systems into a shared framework suitable for multiple downstream tasks. Yang et al. (2023b) fine-tune LLMs to perform recommendation tasks, producing a novel framework for building recommendation systems. Chu et al. (2023) merge different recommendation systems to address the challenge of effectively integrating the commonsense and reasoning abilities of LLMs into recommendation systems. Hu et al. (2024) propose a sequential recommendation framework to preserve fine-grained item textual information.
Zero-Shot Recommendation System.
In addition to the above research, some works focus on zero-shot recommendation systems Wang and Lim (2023); Hou et al. (2024). By utilizing in-context learning, Hou et al. (2024) propose a zero-shot sequential recommendation system. Furthermore, they show that LLMs in recommendation systems have biases due to popularity or item positions in the prompts. Zhang et al. (2023) design recommendation systems with enhanced user-friendliness, allowing users to interact freely with the system and receive more precise recommendations through natural language instructions. For generalizability, Wang et al. (2024f) highlights that current recommendation systems mostly focus on specific tasks and lack the ability to generalize to new tasks. They propose an LLM-powered agent for general recommendation purposes.
Takeaway.
LLM-based recommendation systems use users’ interaction histories to recommend specific items. Several works involve prompting methods, fine-tuning LLMs, or constructing new frameworks. Additionally, Chen (2023) provide a survey of LLM recommendation systems.
3.1.3 Personalized Search
In this section, we introduce applications of LLMs in personalized search systems. Traditional search systems provide a list of relevant search results based on input queries, often limited to simple queries. These results comprise various sources of information, which can be hard to distinguish and organize. Compared to traditional search systems, personalized search systems are based on complex input queries and past interactions to infer user preferences. They synthesize information from multiple sources and present the results in a cohesive, natural language form.
Spatharioti et al. (2023) demonstrate that LLM-based search systems improve users’ performance in certain situations. Ziems et al. (2023) suggest that LLMs act as built-in search engines given few-shot demonstrations. Specifically, LLMs can generate correct web URLs for corresponding documents. Therefore, personalized search has gained popularity as a research field.
There exists a line of investigation focusing on cognitive memory mechanism-based search systems. Zhou et al. (2021b) propose a method aimed at addressing the issue of data sparsity by using a contrastive sampling method to pre-train the ranking model. Furthermore, Zhou et al. (2021a) integrates LLMs with a cognitive memory mechanism, creating a model capable of accessing recent history without processing the entire user history. Building on this, Zhou et al. (2024) presents a strategy to combine a cognitive memory mechanism with LLMs to solve the problem of deep learning-based personalized search, which requires a large amount of training data. This enables LLMs to efficiently retrieve memory.
Empirically, Sharma et al. (2024) conducts experiments to investigate how LLM-powered search systems could lead to opinion polarization. Baek et al. (2024); Salemi and Zamani (2024) leverage search engine results to enhance LLM personalization.
Takeaway.
LLM-based personalized search systems still require further development to ensure that LLMs are efficient and trustworthy with the search results. Li et al. (2024d) provide a comprehensive investigation of LLM-based search and recommendation systems.
3.1.4 Personalized Healthcare
LLMs for the Biomedical Domain.
LLMs have exhibited expert-level capabilities in general biomedical domain tasks, including medical text generation Hendrycks et al. (2021), medical question answering Jin et al. (2020, 2019), and medical document classification Cohan et al. (2020). Multi-modality models for the biomedical domain have the potential to integrate into people’s everyday lives Milne-Ives et al. (2020); Abbasian et al. (2024b). As healthcare assistance varies from person to person, a personalized health wellness assistant possesses a database of users’ preferences and medical history, enabling it to provide precise medical advice or diagnoses.
Personalized LLMs for Healthcare.
Abbasian et al. (2024a) propose openCHA, a framework for personalized healthcare that integrates knowledge from external data sources and personalized health data, and processes it using machine learning modules to address personalized medical problems. Following openCHA, Abbasian et al. (2024c) infuse domain-specific knowledge to effectively utilize health data, knowledge bases, and analytical tools for diabetes-related questions. MaLP Zhang et al. (2024a) combines parameter-efficient fine-tuning Ding et al. (2023) (PEFT) with a memory retrieval module to generate personalized medical responses. Specifically, they separate the memory module into short-term memory (STM), which stores recent knowledge, and long-term memory (LTM), which retains frequently used knowledge from STM.
[…]
Moreover, LLMs also show exceptional potential for psychotherapy Stade et al. (2024); Chen et al. (2023b); Xu et al. (2024). In the near future, personalized LLMs for healthcare could potentially be integrated into clinical scenarios, mental health assessments, and prescribed therapeutic treatments.
Takeaway.
As we witness rapid technological innovation in AI and its broader applications, several questions highlighted in Swift and Allen (2010) need to be considered to design a system that is both reliable and safe. Key issues include clinical decision-making involving AI and humans, and legal considerations concerning the liability associated with these personalized medical assistants.
3.1.5 Personalized Education
LLMs for Education.
LLMs have demonstrated impressive abilities across a broad range of subjects from social sciences to natural sciences, encompassing topics like natural science Wang et al. (2024d), coding Haluptzok et al. (2023), mathematical reasoning Ahn et al. (2024), medicine Saab et al. (2024), finance Yang et al. (2023c), and law Cui et al. (2023). With the depth and breadth of knowledge compressed within LLMs, educators can harness their expertise to facilitate education in several ways. For example, LLMs can provide detailed, step-by-step explanations in the Socratic teaching style (Hao et al., 2024), answer questions on technical and complicated subjects (Arefeen et al., 2023), and automatically summarize lectures to enhance learning experience (Gonzalez et al., 2023). Additionally, LLMs can assist in content creation to help educators develop more diverse and comprehensive materials (Xiao et al., 2023; Heck and Meurers, 2023), and implement automatic grading to ensure consistent grading standards (Matelsky et al., 2023; Tornqvist et al., 2023).
Personalized LLMs for Education.
Recent surveys Kasneci et al. (2023); Gan et al. (2023); Wang et al. (2024c); Jeon and Lee (2023); Huber et al. (2024) have illustrated various opportunities and visions for integrating LLMs into educational environments. These applications range from personalized learning and teaching assistance to homework assessment and feedback. Furthermore, the integration of experts in various subjects to create a unified educator holds tremendous potential for the development of educational LLMs (Li et al., 2023e).
For example, EduChat Dan et al. (2023) pre-trains models on an educational corpus, including textbooks, Chinese poetry, and psychology books, to establish a foundational knowledge base. Then, fine-tune the models on specialized tasks such as essay assessment, Socratic teaching data, emotional support, and retrieval-augmented open-question answering. HumSum Shehata et al. (2023) summarizes personalized lecture transcripts from diverse scenarios, considering factors such as length, depth, tone, and complexity. This is followed by prompt tuning to modify the summary based on the personalization options given by users. Park et al. (2024) incorporate the student’s affective state, cognitive state, and learning style into the prompt to create a personalized conversation-based tutoring system.
Takeaway.
Personalized LLMs have the potential to transform the educational ecosystem into one that is more inclusive and equitable, obviating the need for individuals to pay disproportionate fees. Individuals can access personalized educational content, assist teachers with lecture materials, and receive affordable tutoring, benefitting minority groups, underrepresented individuals, or those with limited access to resources. This concept of personalized LLMs is underscored by Tanya Byron’s statement: "The technology itself is not transformative. It’s the school, the pedagogy, that is transformative." In this context, it is LLMs for personalized education that will bring about the next transformative era.
3.2 LLM Personalized Methods
In this section, we introduce three major methods for personalizing LLMs. Additionally, beyond the categories discussed below, some works are difficult to classify. For example, to achieve customized preference learning, Cheng et al. (2023) propose a three-stage customized reward model learning process. Rocca and Yarkoni (2022) propose a user-encoding method by training with contrastive learning loss. Chu et al. (2023) train a recommendation model by applying a new mask mechanism, span order, positional encoding, and a dynamic position mechanism.
3.2.1 Fine-Tuning Method
Reinforcement Learning from Human Feedback (RLHF; Bai et al., 2022) is the dominant method for aligning LLMs with human intent. Several works have incorporated RLHF for personalizing LLMs. Mondal et al. (2024) propose a method of fine-tuning to create personalized slides from specific research papers. Hwang et al. (2023) use a multi-objective reinforcement learning (MORL) method to personalize robotic agents. Additionally, Jang et al. (2023) show that reframing the task of aligning LLMs with human preferences as a MORL problem can adapt to individual users better, outperforming supervised fine-tuning, RLHF, and prompting-based approaches. Li et al. (2024c) explain the deficit of vanilla RLHF and propose an advanced RLHF algorithm to fine-tune LLM. Kirk et al. (2023) also mention the drawback of RLHF that it may lead to a more polarised view.
In addition to RLHF, some works fine-tune LLMs using different methods: Shea and Yu (2023) propose an offline reinforcement learning method to enhance the personality consistency of dialogue systems. Qin et al. (2023) proposes to use data selection metrics and mini-batch data for more efficient LLM fine-tuning.
3.2.2 Retrieval Augmentation
For retrieval augmentation of personalizing LLMs, Wu et al. (2021) propose an attention-based split memory network. They use a splitting memory network of user profiles and users’ comment histories on Reddit to generate personalized responses. Fu et al. (2022) trained a probabilistic model to select related personal memory as well as knowledge to generate personalized responses. Xu et al. (2022) propose a long-term memory mechanism to continuously extract user persona memory and agent persona memory from dialog history. Zhong et al. (2024) proposed MemoryBank, a long-term memory mechanism focusing on event summary and user portrait. Sun et al. (2024) propose a way to better represent data to make retrieval augmentation more efficient.
Limitations.
Tan et al. (2024) consider that retrieval augmentation might underperform due to unrelated prompts. They propose a fine-tuning way to encode personalized parameters into LLMs. Prior studies have investigated memory-augmented techniques for prompting the LLM with pre-stored user-specific information to generate personalized responses for new queries. However, Zhang et al. (2024b) argues that this approach fails to capture fine-grained details.
3.2.3 Prompting Method
In this section, we introduce the prompting method to personalize LLMs. The prompting method is also the primary way to personalize LLMs. Following Tan et al. (2024), we categorize prompt-based personalization methods fall into three categories.
Vanilla Personalized Prompt.
For vanilla personalized Prompt, they directly prompt user preference and the required task to LLM. Dai et al. (2023) enhance the LLM’s recommendation performance with traditional information retrieval (IR) ranking capabilities, including point-wise, pair-wise, and list-wise ranking. Christakopoulou et al. (2023) propose a user journeys method. LLM improves the recommendation system by reasoning through user activities. Additionally, Zhiyuli et al. (2023) employs a collaborative filtering algorithm to construct a book recommendation system.
Retrieval-augmented Personalized Prompt.
For retrieval-augmented personalized prompts, prompt the LLM to effectively retrieve the relevant content. Mysore et al. (2023) propose a training method for prompt augmentation. Their training method makes LLM able to generate personalized social media posts. Salemi et al. (2024) Offer a comprehensive evaluation framework. Propose two retrieval augmentation approaches. (i) in-prompt augmentation (ii) fusion-in-decoder.i.e., encode personal items separately and integrate them in the decoder. Li et al. (2023a) personalizes LLMs by employing a method inspired by writing education. Specifically, they utilize the following steps to enhance LLMs’ generation ability: multiple stages including retrieval, ranking, summarization, synthesis, and generation.
Profile-Augmented Prompt.
Richardson et al. (2023) notes that incorporating user history data into the prompt to personalize LLMs can lead to inputs exceeding limitations and cost issues. Additionally, retrieval-based methods have the problem of potential information loss. Therefore, they propose a task-aware user summaries method to enable LLMs to achieve a deeper understanding of task profiles. Liu et al. (2023b) concentrates on content-based recommendation and examines both open-source and closed-source LLMs. They fine-tune the model for open-source LLMs and employ the prompting method for closed-source LLMs. Li et al. (2024a) train a prompt rewriter to improve prompt for better-personalized text generation.
4 LLM Personality Evaluation
In the previous sections, we summarize current progress in role-playing and personalized LLMs. Equivalently important is the research about quantifying and evaluating LLMs, like personality traits, and psychological behaviors. Understanding correct evaluation techniques allows us to better align LLM with the assigned role-playing model.
The most straightforward evaluation for LLM role-playing ability is personality. Some work evaluating LLMs with human personality evaluation, like Big Five Jiang et al. (2023); Sorokovikova et al. (2024) and MBTI Pan and Zeng (2023); Song et al. (2024). There are arguments about the suitability to simply "transfer" human psychometric tests to LLMs Dorner et al. (2023). Other works on creating a systematic and quantitative evaluation inventory based on Big Five tailored for LLMs Jiang et al. (2024). Jiang et al. (2023); Sorokovikova et al. (2024) evaluate LLM personality based on the Big Five Personality Inventory (BFI) test [^2] Big Five Personality Factors (John et al., 1999) are Openness, Conscientiousness, Extroversion, Agreeableness, and Neuroticism. and story writing test.
In the BFI evaluation, LLMs often can reflect their assigned persona accurately. Moreover, their persona often influences their linguistic style and personality consistency Frisch and Giulianelli (2024); Jiang et al. (2023). While previous work solely focuses either on semantic accurate output and personality consistent dialogue, Harrison et al. (2019) explore simultaneously controlling semantic accuracy and personality consistency. Jiang et al. (2024) introduce Machine Personality Inventory (MPI) for evaluating LLMs’ personality traits. They use Big Five Personality Factors to evaluate each personality trait consisting of a series of descriptions and a set of options and measure the mean and standard deviation of each trait. By comparing with human evaluation they found out that internal consistency (standard deviation) correlates with model capability.
Pan and Zeng (2023) evaluate LLM with the MBTI test measuring whether LLM possesses a human-like personality. Based on MBTI test results, they conclude that different LLMs possess different MBTI types, and are often affected by their training corpus. Moreover, using simple prompt-tuning have a hard time changing the MBTI type of LLMs. Wang et al. (2024e) evaluate role-playing agents personality fidelity by interviewing character agents for personality tests. Then based on the interview results they ask LLM to rate the score of each personality dimension. They found out that based on their measured personalities it matches well with the assigned character’s personality traits. Song et al. (2024) propose an alternative evaluation method, instead of asking LLM questions and evaluating in Likert scale.
LLM Psychological Evaluation.
In Huang et al. (2023), they recognize personality being only one aspect of psychological evaluation, they propose a framework PsychoBench for evaluating various psychological aspects of LLMs including: personality traits (i.e. Big Five Inventory), interpersonal relationships, motivational tests, emotional abilities. The results of LLMs personality agree with Jiang et al. (2024); Pan and Zeng (2023) that different LLMs exhibit distinct personality traits.
Potential Issues.
Dorner et al. (2023) argue human personality evaluation (e.g. Big Five) cannot simply "transfer" to LLM. They experimentally show that Agree Bias (tendency to agree with both true and false key items, regardless of the actual item content) of LLM like GPT-4, GPT-3.5, and Llama-2 is significantly higher than that observed in human respondents. Other works Song et al. (2024); Jiang et al. (2024) also show that LLMs sometimes do not have consistent personalities like humans. Additionally, Fang et al. (2023) addressed challenges in test-based personality computing research, including personality taxonomies, measurement quality, datasets, performance evaluation, and modeling choice. However, the question of whether it is possible to adopt human personality evaluation or do we need to design alternative test specifically for LLMs remains open.
LLMs as Evaluator.
At the core of LLMs, the evaluator is replacing human or standardized metrics with LLMs as an evaluator. In Rao et al. (2023) they use LLM as a generic MBTI personality evaluator. They construct prompts by combining with original statements from MBTI questions with randomly permute options for unbiased reasons. In PsychoGAT Yang et al. (2024), they use LLM-based gamified assessments to make psychological diagnostics instead of based on standardized scales. The effectiveness of LLMs as personality evaluators shows impressive ability if provided with CoT prompting Ji et al. (2023). On the other end, there are specialized LLMs as evaluators for evaluating other LLMS: PandaLM Wang et al. (2023c) to distinguish the superior model from several LLMs; Auto-J Li et al. (2023b) for evaluating alignment of the models on generality, flexibility, and generality; CritiqueLLM Ke et al. (2023) aim at provide an explainable evaluation. They generate their training data from GPT-4 to obtain referenced and reference-free evaluation data, and fine-tuning CritiqueLLM to generate the critiques.
Takeaway.
To this day, there is still no unified and widely accepted way of precisely quantifying personality in LLMs. Do we evaluate LLMs the same as we evaluate human personality? If not, what are the alternative evaluation standards? What are some social and psychological impacts of different LLMs personalities? This will require further research as it is a relatively new area. As the capacity of LLMs increases beyond our imagination, it is crucial to develop a systematic measure to assess the psychological aspects of these LLMs. Just as personality and psychological traits are important predictors of a wide range of consequential life outcomes for humans, including physical and mental health, career success, social relationships, and major life trajectories Roberts et al. (2007); Ozer and Benet-Martinez (2006); Judge et al. (1999); McCrae and John (1992); Hampson (2012). It is crucial to measure personality and psychological traits in LLMs. This could be vital in the future as they take on more advanced roles and capabilities in society.
5 Challenges and Future Directions
5.1 Safety, Privacy, and Bias Issues
With the rapid development of LLM role-playing and LLM personalization, current studies have highlighted several vital yet often neglected issues that should be taken into consideration. In this section, we outline three issues: safety, privacy, and bias, advocating for the research community to further study these aspects in alignment with human values and to provide more reliable research (Si et al., 2022).
Safety.
Several researchers have explored methods to bypass safety alignment, including CipherChat (Yuan et al., 2023) and Prompt Injection (Greshake et al., 2023). In particular, Chen et al. (2023c) and Li et al. (2024b) demonstrate the negative behaviors in LLM role-playing. Moreover, Jin et al. (2024a) and Shah et al. (2023) successfully manipulate LLMs to collaboratively jailbreak. Another safety issue is toxicity. Deshpande et al. (2023) presents that LLMs consistently exhibit toxicity about a range of topics when assigned personas. All of these works demonstrate the discovery of unsafe problems, indicating a need for more efforts to solve these.
Privacy.
Since LLM personalization heavily relies on users’ personas, including personal information and historical behaviors, privacy is a top priority in this field. Recently,  Wang et al. (2024a) discovered that using the Membership Inference Attack can leak personal information in the training data, raising concerns about encoding personal data into models. Although several research provides methods to address this personal information leakage (Lukas et al., 2023; Gambarelli et al., 2023; Huang et al., 2022; Chen et al., 2023d; Zanella-Béguelin et al., 2020), the research community still needs to put more effort into analyzing privacy issues and developing more acceptable personalized LLMs.
Bias.
While a large number of studies focus on enhancing task performance by LLM role-playing, fewer works explore the disadvantages of assigning personas to LLMs. In this context, Gupta et al. (2023), as one of the first studies, highlights the deep-seated stereotypical biases found in LLMs assigned with socio-demographic personas. However, there is still ample room for investigating and mitigating bias issues in the field.
5.2 Lack of Datasets and Benchmarks
Several tasks in LLM role-playing require specific data formats or environmental information. For instance, web environments require HTML format, while game environments need settings and tools. Additionally, user persona modeling in personalized dialogue lacks contradictory persona datasets, which would more accurately represent real human behaviors (Kim et al., 2024b). Furthermore, LLM personalization faces a scarcity of high-quality personal data for model training due to privacy concerns. Although several personalization methods may exist, researchers are unable to thoroughly evaluate their performance. On the other hand, existing benchmarks for both LLM role-playing and LLM personalization are limited and lack comprehensive evaluations across various dimensions (Chang et al., 2023a). Therefore, expanding datasets and benchmarks for specialized usage and personal information under privacy protection is a potential future direction.
5.3 Multilingual and Multimodal Framework
Recent studies have proposed frameworks that harness the capabilities of LLMs for role-playing and personalization. We suggest extending these works into multilingual and multimodal settings. In terms of the multilingual aspect, although there are some multilingual persona datasets, little research has been conducted to experiment with and develop frameworks using them. As knowledge transformation between languages is vital (Tanwar et al., 2023), multilingual frameworks could be a potential future direction. Regarding the multimodal aspect, several studies have been conducted in web environments leveraging LLMs, where models can utilize both text and visual information (Koh et al., 2024; Liu et al., 2024). However, for other environments and LLM personalization, there is still a lack of these kinds of multimodal frameworks for better utilizing LLM capabilities.
5.4 Connecting to Psychology Research
LLMs have shown tremendous potential, however, whether LLMs pose an existential risk to humankind still requires further investigations. An increasingly popular trend of such evaluation is to draw inspiration from the Psychology field. One of the most comprehensive evaluation methods is PsychoBench (Huang et al., 2023), which comprises four aspects: personality traits, interpersonal relationships, motivational tests, and emotional abilities. Studies by Jiang et al. (2023) and Sorokovikova et al. (2024) present that LLMs exhibit distinct personality traits. Nonetheless, as discussed in Section 4, there is currently no unified and widely accepted method to quantify the evaluation of LLMs. Thus, conducting comprehensive psychological research on the personality of LLMs, or even on persona-assigned LLMs, could be a meaningful direction for future research.
6 Conclusion
Leveraging personas, LLMs can generate tailored responses, effectively adapting to a wide range of scenarios. In this survey paper, we outline two lines of work – role-playing and personalization – for research of personas in the era of LLMs. We first present LLM role-playing, introducing their agentic schemas and emergent behaviors under various interactive environments. We then elaborate on LLM personalization, categorizing different tasks and associated methods for user alignment. Lastly, we highlight current challenges and promising future directions.
To the best of our knowledge, we present the first survey, under the unified view of persona, for LLMs role-playing and personalization. We hope this work serves as an introductory guide for beginners to the field and a stepping stone to foster future endeavors for the community.
Acknowledgements
This work was financially supported by the National Science and Technology Council (NSTC) in Taiwan, under Grants 112-2223-E-002-012-MY5 and 111-$2$-E-002-013-MY3, and from the Featured Area Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (113L900901/113L900902/113L900903).
Appendix A Personalized Dialogue Generation Example
A.1 ToD Modeling Example
In Table 5 we provide an example for Task-oriented Dialogue Modeling Section 3.1.1.
Table 5: Dialogue Example from MultiWOZ Dataset Zang et al. (2020).
Domain Dialogue, Slots, and State
Restaurant [USER:] I need a place to dine in the center that’s expensive.
Slots: restaurant-area: centre, restaurant-pricerange: expensive
State: active_intent: find_restaurant
[SYSTEM:] I have several options for you; do you prefer African, Asian, or British food?
State: active_intent: find_restaurant, requested_slots: restaurant-food
[USER:] Any sort of food would be fine, as long as it is a bit expensive. Could I get the phone number for your recommendation?
Slots: restaurant-area: centre, restaurant-pricerange: expensive, restaurant-food
State: active_intent: find_restaurant, requested_slots: restaurant-phone
[SYSTEM:] There is an African place named Bedouin in the center. How does that sound?
Slots: restaurant-name: Bedouin
[USER:] Sounds good, could I get that phone number? Also, could you recommend me an expensive hotel?
Slots: restaurant-area: centre, restaurant-name: Bedouin, restaurant-pricerange: expensive, restaurant-phone
State: active_intent: find_restaurant, requested_slots: restaurant-phone
Hotel [SYSTEM:] Bedouin’s phone is 01223367660. As far as hotels go, I recommend the University Arms Hotel in the center of town.
Slots: hotel-name: University Arms Hotel
State: active_intent: find_hotel
[USER:] Yes. Can you book it for me?
State: active_intent: find_hotel, requested_slots: hotel-booking
[SYSTEM:] Sure, when would you like that reservation?
[USER:] I want to book it for 2 people and 2 nights starting from Saturday.
Slots: hotel-bookday: Saturday, hotel-bookpeople: 2, hotel-bookstay: 2
State: active_intent: book_hotel, slot_values: hotel-bookday: Saturday, hotel-bookpeople: 2, hotel-bookstay: 2
[SYSTEM:] Your booking was successful. Your reference number is FRGZWQL2. May I help you further?
Slots: hotel-bookday: Saturday, hotel-bookpeople: 2, hotel-bookstay: 2, reference number: FRGZWQL2
[USER:] That is all I need to know. Thanks, goodbye.
[SYSTEM:] Thank you so much for Cambridge TownInfo center. Have a great day!
A.2 User Persona Modeling Example
In Section A.2 we provide and example for Persona Modeling Example Section 3.1.1.
Table 6: User Persona Modeling Example from Persona-Chat Zhang et al. (2018a).
Persona Chat
I fly airplanes. I enjoy building computers. My favorite band is tool. I am in the army. I dropped out of college. [PERSON 1:] Hello how are u tonight
[PERSON 2:] Hi. I am okay. tired, but okay. how are you ?
[PERSON 1:] I am doing good should be sleeping i have school but can’t sleep
[PERSON 2:] I did not finish school, I enlisted in the army instead .
[PERSON 1:] Wow I am only 14 so I can’t do that just yet but I hope too
[PERSON 2:] nice. stay in school and work hard .
[PERSON 1:] I try i like video games and race cars
[PERSON 2:] I like video games too, fallout is my favorite.
[PERSON 1:] I am a call of duty girl i can’t wait for the new one
[PERSON 2:] My younger brother is a cod player too. he is pretty good .
[PERSON 1:] I have three best friends but lots of other friends that play it
[PERSON 2:] I have a best friend, she is a pilot like me.
[PERSON 1:] What kind of plane do u fly
[PERSON 2:] A bomber, it is awesome. do you want to take lessons
[PERSON 1:] I am kinda afraid of heights so not sure flying is for me
[PERSON 2:] You should at least try to go up in a plane, it is a blast.